OpenStudy (noahbred):

ANY ONE HELPP GIVING MEDALS? Andy’s golf club works with 95% efficiency. If Andy puts 120 N-m of work into his swing, how much work will the golf club put out? (Remember to rearrange the formula to find output.) 95 N-m 125 N-m 215 N-m 114 N-m

OpenStudy (anonymous):

do you know the formula?

OpenStudy (noahbred):

Efficiency = Wout (work output)/Win (work input) × 100?

OpenStudy (pawanyadav):

Let x be the output. Then. efficiency=x/input So. 95=(x/120) ×100 Can you find x

OpenStudy (noahbred):

12000

OpenStudy (pawanyadav):

No. 95/100=x/120 So x=?

OpenStudy (noahbred):

0.95?

OpenStudy (pawanyadav):

.95=x/120 Multiply both side by 120 0.95*120=x/120 *120 =>x=0.95*120

OpenStudy (pawanyadav):

Can you multiply them now

OpenStudy (noahbred):

114

OpenStudy (pawanyadav):

Yea. Correct

OpenStudy (anonymous):

well done :)

OpenStudy (noahbred):

thankss

OpenStudy (pawanyadav):

Welcome